# No Dimensional Sampling Coresets for Classification

### v1

@@ Abstract
Keywords. coreset, classification, sensitivity sampling, Radamacher complexity
@@ 1 Introduction
Algorithms for coresets are often described with respect to a finite sample $X\subset\mathcal{X}$ where we then assume $P$ has support limited to and uniform over $X$. We will return later to the integral of $P$ definition to state more general bounds. Then a very common approach to create a coreset is via sensitivity sampling (Feldman and Langberg, 2011). This samples points from $X$ proportional to their sensitivity score $s(x)$. Via an importance sampling argument, one can show the size of the sample $Y$ needed to induce an $\varepsilon$-coreset grows proportionally to the total sensitivity $S=\sum_{x\in X}s(x)$.
For classification, consider a non-increasing function $\phi:\mathbb{R}\longrightarrow(0,\infty)$ and a set $\mathcal{W}\subseteq\mathbb{R}^{d}$. Define $\ell_{\phi}(x,w)=\phi(\langle x,w\rangle)$; this aligns with scenarios like Logistic regression when $\phi(t)=\log(1+e^{-t})$. However, in (Munteanu et al., 2018), it was proven that certain types of non-increasing functions $\phi$, specifically those where $\frac{\phi(x)}{\phi(-x)}$ approaches zero as $x$ goes to infinity, can lead to $\varepsilon$-coresets that encompass all points in $X$ (see Theorem 3.1 in (Tolochinksy et al., 2022) which refines (Munteanu et al., 2018)). To construct a smaller coreset, some additional problem structure needs to be introduced. We mostly focus on the natural way of regularizing the problem (Shalev-Shwartz and Ben-David, 2014), initiated in the context of coresets by Tolochinksy et al. (2022). For a positive integer $k$, define $\ell_{k,\phi}(x,w)=\phi(\langle x,w\rangle)+\frac{1}{k}\|w\|^{2}_{2}$. An $\varepsilon$-coreset for $(X,p,\mathcal{W},\ell_{k,\phi})$ was termed a coreset for a monotonic function $\phi$ (Tolochinksy et al., 2022; Curtin et al., 2020). Others approached this issue in varied ways. For example, Munteanu et al. (2018) introduced a complexity measure $\mu(X)$ that quantifies the hardness of compressing a dataset for logistic regression, and showed sublinear-sized coresets can be founding depending on this parameter. Their results were polynomially improved by Mai et al. (2021) in terms of $\mu(X)$ and $d$. Other work (Mirzasoleiman et al., 2020) considers greedily selected coresets for learning classifiers under incremental gradient descent where regularization is again needed, but now to ensure strong convexity.
@@
  • ReLU function: $\text{ReLU}(t)=\max(0,t)$
@@
  • We provide a new general sensitivity sampling bound based on Radamacher complexity (Theorem 2), which leads to these no dimensional bounds, and we believe will be of general interest.
@@
$\operatorname{{\rm logistic}}$ bounded $\mu(X)$ sqrt lev. scores $\begin{array}[]{c}O^{*}\left(\frac{d^{3}\mu^{3}(X)}{\varepsilon^{4}}\right)\\ O^{*}\left(\frac{\sqrt{n}d^{\frac{3}{2}}\mu(X)}{\varepsilon^{2}}\right)\end{array}$ Munteanu et al. 2018
$\operatorname{{\rm logistic}},\operatorname{{\rm svm}},\text{ReLU}$ bounded $\mu(X)$ $\ell_{l}$ Lewis $\begin{array}[]{c}O\left(\frac{\mu^{2}(X)d\log(d/\varepsilon)}{\varepsilon^{2}}\right)\end{array}$ Mai et al. 2021
$\operatorname{{\rm logistic}},\operatorname{{\rm svm}}$ $\begin{array}[]{c}\|x_{i}\|_{2}\leq 1\ \forall i\\ \text{reg. term: }\frac{1}{k}\|w\|_{2}^{2}\\ \|w\|_{2}\leq 1\end{array}$ sensitivity $\begin{array}[]{c}O^{*}\left(\frac{kd^{2}\log n\log k}{\varepsilon^{2}}\right)\end{array}$ Tolochinksy et al. 2022
@@
$\operatorname{{\rm logistic}},\operatorname{{\rm svm}}$ $\begin{array}[]{c}\|x_{i}\|_{2}\leq 1\ \forall i\\ \text{reg. term: }\frac{1}{k}\|w\|_{2}^{2}\\[2.84544pt] \frac{1}{k}\|w\|_{2},\text{ or }\frac{1}{k}\|w\|_{1}\end{array}$ uniform $\begin{array}[]{c}O\left(\frac{dk\log k}{\varepsilon^{2}}\right)\end{array}$ Curtin et al. 2020
@@
$\operatorname{{\rm logistic}}$ $\begin{array}[]{c}\|x_{i}\|_{2}\leq 1\ \forall i\\ \text{reg. term: }\frac{1}{k}\|w\|_{2}^{2}\end{array}$ uniform $\begin{array}[]{c}O\left(\frac{dk\sqrt{\log k}}{\varepsilon^{2}}\right)\\[5.69046pt] O\left(\frac{k^{3}}{\varepsilon^{2}\sqrt{\log k}}\right)\end{array}$ Theorems 2825
@@ 1.1 Preliminaries
As a reminder, “for almost all $x\in\mathcal{X}$” signifies that the statement holds for all $x\in\mathcal{X}$ except for a subset of measure zero with respect to $P$. Given that a subset of measure zero does not affect sampling or integral outcomes, without loss of generality, we can assume the inequality holds for all $x\in\mathcal{X}$. It can be done by either setting the value of each $f\in\mathcal{F}$ to zero for every $x$ in that measure zero subset or by removing that measure zero subset from $\mathcal{X}$. We maintain this assumption going forward. The value $S=\int s(x){\rm d}P(x)$ will be referred to as the total sensitivity for $(\mathcal{X},P,\mathcal{F})$ with respect to the function $s(\cdot)$. The sensitivity normalized probability measure for $(\mathcal{X},P,\mathcal{F})$ is defined as ${\rm d}Q=\frac{s(x)}{S}{\rm d}P$ or equivalently $Q(A)=\int_{A}\frac{s(x)}{S}{\rm d}P(x)$ for each $P$-measurable $A$. Given that $S=\int s(x){\rm d}P(x)$, it follows that $Q$ represents a probability measure over $\mathcal{X}$ since $\int_{\mathcal{X}}{\rm d}Q(x)=\int_{\mathcal{X}}\frac{s(x)}{S}{\rm d}P(x)=1$. An $s$-sensitivity sample from $\mathcal{X}$ with a size of $m$ is a set of $m$ iid draws $x_{1}\ldots,x_{m}$ from $\mathcal{X}$ based on $Q$ which will be expressed as as $x_{1:m}\sim Q$. To simplify, one can imagine iid sampling from $\mathcal{X}$ using the probability distribution $q(x)=\frac{s(x)}{S}p(x)$. Define the $s$-augmented family of $(\mathcal{X},P,\mathcal{F})$ as
@@
$\Big{|}\int_{\mathcal{X}}f(x){\rm d}P(x)-\sum_{i=1}^{m}\frac{Sf(x_{i})}{ms(x_{i})}\Big{|}\leq\varepsilon\int_{\mathcal{X}}f(x){\rm d}P(x)\Longleftrightarrow\quad\left|1-\frac{1}{m}\sum_{i=1}^{m}T_{f}(x_{i})\right|\leq\varepsilon.$ (3)
@@
Remark 1
We assume, without explicitly stating, that the supremum over the family $\mathcal{F}$ in this definition is measurable. This assumption is consistently applied throughout this paper for other suprema over a class of functions. It is worth noting that while this assumption may not be applicable to arbitrary function classes, it holds true for the hypothesis sets commonly encountered in practical machine learning scenarios. For example, when either $\mathcal{X}$ or $\mathcal{F}$ is countable, this assumption remains valid. Furthermore, all the specific cases addressed in this work meet this requirement.
@@ 1.2 Our Contributions
$\Big{|}\int f(x){\rm d}P(x)-\sum_{i=1}^{m}\frac{Sf(x_{i})}{ms(x_{i})}\Big{|}(2\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{F}})+t)\int f(x){\rm d}P(x)\quad\forall f\in\mathcal{F}.$
This theorem, which is pivotal in our new results, is an adaptation of the central result from (Feldman and Langberg, 2011; Braverman et al., 2021; Feldman et al., 2020), extensively employed in coreset construction results, but now centered on Rademacher Complexity instead of VC-dimension.
@@
Theorem 3
Let $(\mathcal{X},P,\mathcal{F})$ be a positive definite tuple, and $s\colon\mathcal{X}\longrightarrow(0,\infty)$ be an upper sensitivity function for it with the total sensitivity $S=\int s(x){\rm d}P(x)$. There is a universal constant $C$ such that, for any $10<\varepsilon,\delta<$ 1, if $m\geq\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, where $d={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{F}},\succ))$, then, with probability at least $1-\delta$, any $s$-sensitivity sample $x_{1}\ldots,x_{m}$ from $\mathcal{X}$ satisfies
$\Big{|}\int f(x){\rm d}P(x)-\frac{1}{m}\sum_{i=1}^{m}\frac{Sf(x_{i})}{s(x_{i})}\Big{|}\leq\varepsilon\int f(x){\rm d}P(x)\quad\quad\forall f\in\mathcal{F}.$
This theorem, with variants previously shown in (Braverman et al., 2021; Feldman et al., 2020), are usually constrained by the assumption that $\mathcal{X}$ is finite, which has been extensively used to construct coresets. However, in this work we remove this restriction. Section 5 showcases a novel and simple proof for this theorem, leveraging Fubini’s Theorem (Fubini, 1907) [^1] Fubini’s Theorem is applicable in our context, given we are dealing with $\sigma$-finite measures.. Moreover, in instances where Rademacher complexity refines VC-dimension, our new result Theorem 2 provides improved bounds. Next we describe several applications stemming from these two theorems.
@@
Theorem 5
Let $P$ be a well-behaved probability measure over $\mathbb{R}^{d}$. If $s(\cdot):\mathbb{R}^{d}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}_{\phi,k}$ with total sensitivity $S$ and $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$, then any $s$-sensitivity sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{S}{ms(x_{i})}$ provides an $\varepsilon$-coreset for $(\mathbb{R}^{d},\ P,\ \mathbb{R}^{d},\ \ell_{\phi,k})$ with probability at least $1-\delta$, where $C=\left(2LE_{1}+\phi(0)\right)\max\left(2E_{1}k,\frac{1}{E_{2}}\right)+8LkE_{1}+1$.
When $E_{1}$ is independent of dimension, such as bounded or normalized data, this theorem gives an $\varepsilon$-coreset for $(\mathbb{R}^{d},\ P,\ \mathbb{R}^{d},\ \ell_{\phi,k})$ with size independent of dimension $d$ and input size $n$. In contrast, the coresets for (Tolochinksy et al., 2022), also considering bounded data, used VC-dimension (via Theorem 3), so their size depends on $d$ and $n$.
@@
It should be emphasized that the last two results in Theorem 6 was successfully improved to $O(\frac{k}{\varepsilon^{2}}(d\log k+\log\frac{1}{\delta})$ for $\mathcal{W}=\mathbb{R}^{d}$ and $k=\Theta(n^{1-\kappa})$ for some $\kappa\in(0,1)$ by Curtin et al. (2020). The strategy utilized by Tolochinksy et al. (2022) and Curtin et al. (2020) to demonstrate this theorem can be encapsulated as: (1) find an upper bound for VC-dimension, (2) approximate a sensitivity upper bound and thus the total sensitivity, (3) sample $m$ points form $X$ according to an appropriate distribution, and (4) use Theorem 3 to complete the argument. The argument by Tolochinksy et al. (2022), however, encounters a critical issue. To apply Theorem 3, we must replace $d$ with the VC-dimension of the linked-range space of the $s$-augmented family of $(\mathcal{X},P,\mathcal{L}_{\phi,k})$, i.e. ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\phi,k}},\succ))$, while they directly used the VC-dimension of linked-range space of $\mathcal{F}$ (instead of $\mathcal{T}_{\mathcal{L}_{\phi,k}}$). Our work addresses this issue in that it leads to sensitivity functions for which we can compute the VC-dimension of the augmented function space. Furthermore, by instead bounding Radamacher complexity in step (1), we can then use our new Theorem 2 in step (4) to obtain new bounds. These are summarized in Table 2 and the next theorem; details are in appendix.
Theorem 7
@@
$\sigma$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E^{2}_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Theorem 24
$\sigma$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E^{2}_{1}$ $O(E_{1}^{2}k)$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Theorem 25
$\operatorname{{\rm logistic}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(\frac{A^{2}k}{\sqrt{\log(A^{2}k)}})$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Theorem 28
$\operatorname{{\rm logistic}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(\frac{A^{2}k}{\sqrt{\log(A^{2}k)}})$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Theorem 25
$\operatorname{{\rm logistic}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(\frac{E_{1}^{2}k}{\sqrt{\log(E_{1}^{2}k)}})$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Theorem 31
$\operatorname{{\rm logistic}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(\frac{E_{1}^{2}k}{\sqrt{\log(E_{1}^{2}k)}})$ $O(1)$ $\frac{s}{S}p$ $\frac{CS}{\varepsilon^{2}}\left(d^{2}\log S+\log\frac{1}{\delta}\right)$ Theorem 33
$\operatorname{{\rm svm}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(A^{2}k)$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Theorem 36
$\operatorname{{\rm svm}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(A^{2}k)$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Theorem 38
$\operatorname{{\rm svm}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Theorem 40
$\operatorname{{\rm svm}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(E_{1}^{2}k)$ $O(1)$ $\frac{s}{S}p$ $\frac{CS}{\varepsilon^{2}}\left(d^{2}\log S+\log\frac{1}{\delta}\right)$ Theorem 41
ReLU $P\left(\|x\|_{2}\geq A\right)=0$ $O((1+A)^{2}k)$ $O((1+A)^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Section 4.4
ReLU $P\left(\|x\|_{2}\geq A\right)=0$ $O((1+A)^{2}k)$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Section 4.4
ReLU $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O((1+E_{1})^{2}k)$ $O((1+E_{1})^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Section 4.4
ReLU $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O((1+E_{1})^{2}k)$ $O(1)$ $\frac{s}{S}p$ $\frac{CS}{\varepsilon^{2}}\left(d^{2}\log S+\log\frac{1}{\delta}\right)$ Section 4.4

### v2

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Algorithms for coresets are often described with respect to a finite sample $X\subset\mathcal{X}$ where we then assume $P$ has support limited to and uniform over $X$. We will return later to the integral of $P$ definition to state more general bounds. Then a very common approach to create a coreset is via sensitivity sampling (Feldman & Langberg, 2011). This samples points from $X$ proportional to their sensitivity score $s(x)$. Via an importance sampling argument, one can show the size of the sample $Y$ needed to induce an $\varepsilon$-coreset grows proportionally to the total sensitivity $S=\sum_{x\in X}s(x)$.
For classification, consider a non-increasing function $\phi:\mathbb{R}\longrightarrow(0,\infty)$ and a set $\mathcal{W}\subseteq\mathbb{R}^{d}$. Define $\ell_{\phi}(x,w)=\phi(\langle x,w\rangle)$; this aligns with scenarios like Logistic regression when $\phi(t)=\log(1+e^{-t})$. However, in (Munteanu et al., 2018), it was proven that certain types of non-increasing functions $\phi$, specifically those where $\frac{\phi(x)}{\phi(-x)}$ approaches zero as $x$ goes to infinity, can lead to $\varepsilon$-coresets that encompass all points in $X$ (see Theorem 3.1 in (Tolochinksy et al., 2022) which refines (Munteanu et al., 2018)). To construct a smaller coreset, some additional problem structure needs to be introduced. We mostly focus on the natural way of regularizing the problem (Shalev-Shwartz & Ben-David, 2014), initiated in the context of coresets by Tolochinksy et al. (2022). For a positive integer $k$, define $\ell_{k,\phi}(x,w)=\phi(\langle x,w\rangle)+\frac{1}{k}\|w\|^{2}_{2}$. An $\varepsilon$-coreset for $(X,p,\mathcal{W},\ell_{k,\phi})$ was termed a coreset for a monotonic function $\phi$ (Tolochinksy et al., 2022; Curtin et al., 2020). Others approached this issue in varied ways. For example, Munteanu et al. (2018) introduced a complexity measure $\mu(X)$ that quantifies the hardness of compressing a dataset for logistic regression, and showed sublinear-sized coresets can be founding depending on this parameter. Their results were polynomially improved by Mai et al. (2021) in terms of $\mu(X)$ and $d$. Other work (Mirzasoleiman et al., 2020) considers greedily selected coresets for learning classifiers under incremental gradient descent where regularization is again needed, but now to ensure strong convexity.
@@
  • ReLU function: $\operatorname{{\rm ReLU}}(t)=\max(0,t)$
@@
  • We provide a new general sensitivity sampling bound based on Radamacher complexity (Theorem 1.1), which leads to these no dimensional bounds, and we believe will be of general interest.
@@
$\operatorname{{\rm logistic}}$ bounded $\mu(X)$ sqrt lev. scores $\begin{array}[]{c}O^{*}\left(\frac{d^{3}\mu^{3}(X)}{\varepsilon^{4}}\right)\\ O^{*}\left(\frac{\sqrt{n}d^{\frac{3}{2}}\mu(X)}{\varepsilon^{2}}\right)\end{array}$ (Munteanu et al., 2018)
$\operatorname{{\rm logistic}},\operatorname{{\rm svm}},\operatorname{{\rm ReLU}}$ bounded $\mu(X)$ $\ell_{l}$ Lewis $\begin{array}[]{c}O\left(\frac{\mu^{2}(X)d\log(d/\varepsilon)}{\varepsilon^{2}}\right)\end{array}$ (Mai et al., 2021)
$\operatorname{{\rm logistic}},\operatorname{{\rm svm}}$ $\begin{array}[]{c}\|x_{i}\|_{2}\leq 1\ \forall i\\ \text{reg. term: }\frac{1}{k}\|w\|_{2}^{2}\\ \|w\|_{2}\leq 1\end{array}$ sensitivity $\begin{array}[]{c}O^{*}\left(\frac{kd^{2}\log n\log k}{\varepsilon^{2}}\right)\end{array}$ (Tolochinksy et al., 2022)
@@
$\operatorname{{\rm logistic}},\operatorname{{\rm svm}}$ $\begin{array}[]{c}\|x_{i}\|_{2}\leq 1\ \forall i\\ \text{reg. term: }\frac{1}{k}\|w\|_{2}^{2}\\[2.84544pt] \frac{1}{k}\|w\|_{2},\text{ or }\frac{1}{k}\|w\|_{1}\end{array}$ uniform $\begin{array}[]{c}O\left(\frac{dk\log k}{\varepsilon^{2}}\right)\end{array}$ (Curtin et al., 2020)
@@
$\operatorname{{\rm logistic}}$ $\begin{array}[]{c}\|x_{i}\|_{2}\leq 1\ \forall i\\ \text{reg. term: }\frac{1}{k}\|w\|_{2}^{2}\end{array}$ uniform $\begin{array}[]{c}O\left(\frac{dk\sqrt{\log k}}{\varepsilon^{2}}\right)\\[5.69046pt] O\left(\frac{k^{3}}{\varepsilon^{2}\sqrt{\log k}}\right)\end{array}$ Theorems 2.72.8
@@ 1.1 Preliminaries
The value $S=\int s(x){\rm d}P(x)$ will be referred to as the total sensitivity for $(\mathcal{X},P,\mathcal{F})$ with respect to the function $s(\cdot)$. The sensitivity normalized probability measure for $(\mathcal{X},P,\mathcal{F})$ is defined as ${\rm d}Q=\frac{s(x)}{S}{\rm d}P$ or equivalently $Q(A)=\int_{A}\frac{s(x)}{S}{\rm d}P(x)$ for each $P$-measurable $A$. Given that $S=\int s(x){\rm d}P(x)$, it follows that $Q$ represents a probability measure over $\mathcal{X}$ since $\int_{\mathcal{X}}{\rm d}Q(x)=\int_{\mathcal{X}}\frac{s(x)}{S}{\rm d}P(x)=1$. An $s$-sensitivity sample from $\mathcal{X}$ with a size of $m$ is a set of $m$ iid draws $x_{1}\ldots,x_{m}$ from $\mathcal{X}$ based on $Q$ which will be expressed as as $x_{1:m}\sim Q$. To simplify, one can imagine iid sampling from $\mathcal{X}$ using the probability distribution $q(x)=\frac{s(x)}{S}p(x)$. Define the $s$-augmented family of $(\mathcal{X},P,\mathcal{F})$ as
@@
$\Big{|}\int_{\mathcal{X}}f(x){\rm d}P(x)-$ $\sum_{i=1}^{m}\frac{Sf(x_{i})}{ms(x_{i})}\Big{|}\leq\varepsilon\int_{\mathcal{X}}f(x){\rm d}P(x)$ (3)
$\Longleftrightarrow$ $\left|1-\frac{1}{m}\sum_{i=1}^{m}T_{f}(x_{i})\right|\leq\varepsilon.$
@@ 1.2 Our Contributions
$\Big{|}\int f(x){\rm d}P(x)$ $-\sum_{i=1}^{m}\frac{Sf(x_{i})}{ms(x_{i})}\Big{|}$
$\leq$ $(2\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{F}})+t)\int f(x){\rm d}P(x)\quad\forall f\in\mathcal{F}.$
This theorem, which is pivotal in our new results, is an adaptation of the central result from (Feldman & Langberg, 2011; Braverman et al., 2021; Feldman et al., 2020), extensively employed in coreset construction results, but now centered on Rademacher Complexity instead of VC-dimension.
@@
Theorem 1.2.
Let $(\mathcal{X},P,\mathcal{F})$ be a positive definite tuple, and $s\colon\mathcal{X}\longrightarrow(0,\infty)$ be an upper sensitivity function for it with the total sensitivity $S=\int s(x){\rm d}P(x)$. There is a universal constant $C$ such that, for any $10<\varepsilon,\delta<$ 1, if $m\geq\frac{CS}{\varepsilon^{2}}\left(\mathsf{VC}\log S+\log\frac{1}{\delta}\right))$, where $\mathsf{VC}={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{F}},\succ))$, then, with probability at least $1-\delta$, any $s$-sensitivity sample $x_{1}\ldots,x_{m}$ from $\mathcal{X}$ satisfies
$\Big{|}\int f(x){\rm d}P(x)$ $-\frac{1}{m}\sum_{i=1}^{m}\frac{Sf(x_{i})}{s(x_{i})}\Big{|}$
$\leq$ $\varepsilon\int f(x){\rm d}P(x)\quad\quad\forall f\in\mathcal{F}.$
This theorem, with variants previously shown in (Braverman et al., 2021; Feldman et al., 2020), are usually constrained by the assumption that $\mathcal{X}$ is finite, which has been extensively used to construct coresets. However, in this work we remove this restriction. This general case was also handled in (Langberg & Schulman, 2010). Section 2.2 showcases a novel and simple proof for this theorem, leveraging Fubini’s Theorem (Fubini, 1907) [^1] Fubini’s Theorem is applicable in our context, given we are dealing with $\sigma$-finite measures.. Moreover, in instances where Rademacher complexity refines VC-dimension, our new result Theorem 1.1 provides improved bounds. Next we describe several applications stemming from these two theorems.
@@
Theorem 1.4.
Let $P$ be a well-behaved probability measure over $\mathbb{R}^{d}$ and $\phi$ be an $L$-Lipschitz function. If $s(\cdot):\mathbb{R}^{d}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}_{\phi,k}$ with total sensitivity $S$ and $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$, then any $s$-sensitivity sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{S}{ms(x_{i})}$ provides an $\varepsilon$-coreset for $(\mathbb{R}^{d},\ P,\ \mathbb{R}^{d},\ \ell_{\phi,k})$ with probability at least $1-\delta$, where $C=\left(2LE_{1}+\phi(0)\right)\max\left(2E_{1}k,\frac{1}{E_{2}}\right)+8LkE_{1}+1$.
When $E_{1}$ is independent of dimension, such as bounded or normalized data, this theorem gives an $\varepsilon$-coreset for $(\mathbb{R}^{d},\ P,\ \mathbb{R}^{d},\ \ell_{\phi,k})$ with size independent of dimension $d$ and input size $n$. In contrast, the coresets for (Tolochinksy et al., 2022), also considering bounded data, used VC-dimension (via Theorem 1.2), so their size depends on $d$ and $n$.
@@
It should be emphasized that the last two results in Theorem 1.5 was successfully improved to $O(\frac{k}{\varepsilon^{2}}(d\log k+\log\frac{1}{\delta})$ for $\mathcal{W}=\mathbb{R}^{d}$ and $k=\Theta(n^{1-\kappa})$ for some $\kappa\in(0,1)$ by Curtin et al. (2020). The strategy utilized by Tolochinksy et al. (2022) and Curtin et al. (2020) to demonstrate this theorem can be encapsulated as: (1) find an upper bound for VC-dimension, (2) approximate a sensitivity upper bound and thus the total sensitivity, (3) sample $m$ points form $X$ according to an appropriate distribution, and (4) use Theorem 1.2 to complete the argument. The argument by Tolochinksy et al. (2022), however, encounters a critical issue. To apply Theorem 1.2, we must replace $d$ with the VC-dimension of the linked-range space of the $s$-augmented family of $(\mathcal{X},P,\mathcal{L}_{\phi,k})$, i.e. ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\phi,k}},\succ))$, while they directly used the VC-dimension of linked-range space of $\mathcal{F}$ (instead of $\mathcal{T}_{\mathcal{L}_{\phi,k}}$). Our work addresses this issue in that it leads to sensitivity functions for which we can compute the VC-dimension of the augmented function space. Furthermore, by instead bounding Radamacher complexity in step (1), we can then use our new Theorem 1.1 in step (4) to obtain new bounds. These are summarized in Table 2 and the next theorem; details are in appendix.
$\sigma$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E^{2}_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Thm D.6
$\sigma$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E^{2}_{1}$ $O(E_{1}^{2}k)$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Thm D.7
$\operatorname{{\rm logistic}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(\frac{A^{2}k}{\sqrt{\log(A^{2}k)}})$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Thm D.10
$\operatorname{{\rm logistic}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(\frac{A^{2}k}{\sqrt{\log(A^{2}k)}})$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Thm 2.8
$\operatorname{{\rm logistic}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(\frac{E_{1}^{2}k}{\sqrt{\log(E_{1}^{2}k)}})$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Thm D.13
$\operatorname{{\rm logistic}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(\frac{E_{1}^{2}k}{\sqrt{\log(E_{1}^{2}k)}})$ $O(1)$ $\frac{s}{S}p$ $\frac{CS}{\varepsilon^{2}}\left(d^{2}\log S+\log\frac{1}{\delta}\right)$ Thm D.15
$\operatorname{{\rm svm}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(A^{2}k)$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Thm D.18
$\operatorname{{\rm svm}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O(A^{2}k)$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Thm D.20
$\operatorname{{\rm svm}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Thm D.22
$\operatorname{{\rm svm}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O(E_{1}^{2}k)$ $O(1)$ $\frac{s}{S}p$ $\frac{CS}{\varepsilon^{2}}\left(d^{2}\log S+\log\frac{1}{\delta}\right)$ Thm D.23
$\operatorname{{\rm ReLU}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O((1+A)^{2}k)$ $O((1+A)^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Sec. D.4
$\operatorname{{\rm ReLU}}$ $P\left(\|x\|_{2}\geq A\right)=0$ $O((1+A)^{2}k)$ $O(1)$ $p$ $\frac{CS}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)$ Sec. D.4
$\operatorname{{\rm ReLU}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O((1+E_{1})^{2}k)$ $O((1+E_{1})^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Sec. D.4
$\operatorname{{\rm ReLU}}$ $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{2}^{2}\leq E_{1}$ $O((1+E_{1})^{2}k)$ $O(1)$ $\frac{s}{S}p$ $\frac{CS}{\varepsilon^{2}}\left(d^{2}\log S+\log\frac{1}{\delta}\right)$ Sec. D.4
Theorem 1.6.